# Velocitune: A Velocity-based Dynamic Domain Reweighting Method for Continual Pre-training

### v1

@@ Abstract
It is well-known that a diverse corpus is critical for training large language models, which are typically constructed from a mixture of various domains. In general, previous efforts resort to sampling training data from different domains with static proportions, as well as adjusting data proportions during training. However, few methods have addressed the complexities of domain-adaptive continual pre-training. To fill this gap, we propose Velocitune, a novel framework dynamically assesses learning velocity and adjusts data proportions accordingly, favoring slower-learning domains while shunning faster-learning ones, which is guided by a scaling law to indicate the desired learning goal for each domain with less associated cost. To evaluate the effectiveness of Velocitune, we conduct experiments in a reasoning-focused dataset with CodeLlama, as well as in a corpus specialised for system command generation with Llama3 and Mistral. Velocitune achieves performance gains in both math and code reasoning tasks and command-line generation benchmarks. Further analysis reveals that key factors driving Velocitune’s effectiveness include target loss prediction and data ordering.
@@ 1 Introduction
Datasets used for pre-training language models (LMs) are typically composed of texts of various meta-attributes like source and focus, referred to as different domains (Du et al., 2022; Azerbayev et al., 2023; Computer, 2023). The distinct characteristics of data from these varying domains, such as focus, quality and quantity, affect the downstream performance of LMs differently (Rozière et al., 2024; Li et al., 2024). Consequently, numerous studies have explored the optimal combination of data from multiple domains to enhance LM performance.
@@
Llama3 (AI@Meta, 2024), GLaM (Du et al., 2022), and Lemma (Azerbayev et al., 2023) employ heuristic methods to iteratively test different ratios by training multiple proxy models and selecting the mixture that demonstrates the best downstream performance. However, these heuristic approaches demand costly large-scale experiments for effective exploration. As a result, recent research focuses on learning an optimal ratio by dynamically adjusting weights during proxy model training (Xie et al., 2024; Fan et al., 2023). For example, Doremi (Xie et al., 2024) implements a method in which a small reference model is initially trained, followed by training a small proxy model using group distributionally robust optimization (Group DRO) (Sagawa et al., 2019) to obtain domain weights.
@@
Domain-adaptive continual pre-training, while sharing some similarities with from-scratch training, presents unique challenges that limit the effectiveness of existing domain reweighting methods. Many existing methods utilise small proxy models to estimate optimal domain weights, which are subsequently transferred to a larger model (Xie et al., 2024; Fan et al., 2023; Azerbayev et al., 2023). However, this approach poses challenges in domain-adaptive continual pre-training, as smaller versions of the base model often do not exist, making it difficult to estimate weights from these proxy models. Another challenge is how to leverage the learning status. Previous methods rely on the distance between the model’s current loss and a target domain loss (Xie et al., 2024; Xia et al., 2023). Nevertheless, this distance-based approach can result in uneven domain emphasis, as domains with larger loss disparities may disproportionately influence learning. This can exacerbate imbalances across domains.
@@
To address these issues, we introduce a novel framework, Velocitune, centered on the concept of learning velocity. In contrast to previous approaches that leverage the distance between the current loss and a target loss, Velocitune more accurately capture how effectively models learn across each domain by establishing learning velocity. During training, domains exhibiting slower learning velocities are given increased weights, while those with faster velocities receive reduced weights, promoting equilibrium in learning progress. To quantify learning velocity, it is crucial to determine both the initial loss, reflecting the model’s already learnt expertise, and the target losses, indicating the desired learning goal for each domain. For a more cost-effective estimation of target losses, we adopt the Chinchilla scaling law (Hoffmann et al., 2022). Instead of using a small model, we predict target reference losses by leveraging models trained on sub-sampled data.
@@
We evaluate the performance of Velocitune in two settings: continual pre-training CodeLlama 7B (Rozière et al., 2024) on a math and coding reasoning dataset as well as Llama3 and Mistral (Jiang et al., 2023) on a system command knowledge corpus. Velocitune demonstrates an average improvement of 1.6% across seven math tasks and 3.8% across two coding tasks compared to baseline models trained with default weights. In addition, Velocitune outperforms the baselines on Llama3, showing improvements of 5.3% and 3.1%, and on Mistral, with gains of 4.4% and 2.6% on the CmdGen-Nvidia and CmdGen-AMD tasks, respectively. Further more, we conduct an in-depth ablation study to identify key factors contributing to the observed improvements.Our findings indicate that, beyond the contribution of reweighted data ratios, the sequence of data ordering might play a significant role in the effectiveness of Velocitune. The results also show that the predicted target loss are critical for the success of the Velocitune. The contributions of this work can be summarised as follows:
@@
Introducing Velocitune, a novel framework designed to adjust data ratios in continual pre-training. It establishes learning velocity to more accurately assess the extent of models’ learning in domains and leverages a scaling law to minimise the costs.
@@
Providing a detailed analysis revealing that data ordering, reweighted data ratios, and predicted target loss significantly contribute to Velocitune’s effectiveness.
@@ 2 Domain-adaptive continual pre-training with Velocitune
In this section, we present a detailed explanation of Velocitune. Our approach focuses on adjusting domain weights based on learning velocity. To quantify it, it is essential to determine both the initial and target losses for each domain. To achieve a cost-effective estimation of target losses, we adopt the Chinchilla scaling law to predict it using models trained using sub-sampled data. During training, we periodically assess learning velocity, increasing the sampling weights of slower domains while reducing weights of faster ones. This adjustment ensures more balanced learning across domains. The methodology is detailed in § 2.1.
@@ 2.1 The Velocitune algorithm
Consider training a language model (LM) on a dataset $S$ consisting of $n$ distinct domains, denoted as $D^{1},D^{2},\dots,D^{n}$. Let ${\mathbb{S}}={S^{1},S^{2},\dots,S^{n}}$ represent the subsets of data corresponding to each domain, where each $S^{i}$ contains both training and evaluation data, denoted as $S_{\text{train},i}$ and $S_{\text{eval},i}$, respectively. The domain weight $w\in\Delta_{n}$ represents the sampling weight assigned to domains.
@@
We first compute the initial loss ${\bm{L}}_{\text{init}}$. This involves evaluating the LM on the full evaluation set ${\mathbb{S}}_{\text{eval}}$, providing an accurate measure of the model’s learnt expertise in each domain. The target loss for each domain is derived using the scaling law (Kaplan et al., 2020). Specifically, in the absence of models with varying sizes, we apply the Chinchilla scaling law by fisrt defining the total dataset size and then training the model on a subsample using the default domain weights, which correspond to the ratio of tokens in each domain. Throughout the training process, multiple checkpoints are saved, and the evaluation losses from these checkpoints are used to fit the scaling law parameters. Finally, we take the fitted function to predict the loss the model could reach for using the entire dataset. More detailed analysis of prediction errors is provided in Appendix B.
@@
After obtaining the initial losses and target losses in each domain, our objective is to iteratively update domain weights, denoted by ${\bm{w}}_{t}$, to minimize the re-weighted training loss effectively. The optimization problem is formulated as follows, encompassing two sequential steps:
(1). Minimise the weighted training loss:
$\min_{\theta}\sum_{t=1}^{T}\sum_{i=1}^{k}w_{t}[i]\cdot\ell_{t}^{i}(\theta)$
@@
where $\ell_{t}^{i}(\theta)$ denotes the training loss at step $t$ for domain $D_{i}$ with model parameters $\theta$. The $w_{t}[i]$ refers to the normalized weight assigned to domain $D_{i}$ at step $t$.
(2). Minimise the weighted sum of learning velocities:
$\min_{w\in\Delta^{k}}\sum_{i=1}^{k}w_{t}[i]\cdot V_{t}[i],$
@@
The first step is to minimise the language modeling Negative Log Likelihood(NLL) loss given the domain weights. The second step focuses on dynamic adjustment such that the weight in the slowest learning velocity domain is maximised. By ensuring a focus on the most challenging scenarios, this helps in improving overall balance during training. Unlike the Group DRO optimization in Doremi (Xie et al., 2024), which updates domain weights based on the loss difference of each data point between a proxy model and a trained reference model, our method does not rely on a fully trained reference model. Consequently, the loss of each data point on the trained model is unobtainable. Instead, after a set number of training steps, we update the domain weights by first calculating the learning velocity, which reflects how quickly the model is learning in a given domain. The learning velocity is defined as:
$V_{t}[i]=\frac{\ell_{t}^{i}(\theta,S^{\prime}_{\text{eval},i})-\ell_{\text{target}}^{i}}{\ell_{\text{init}}^{i}-\ell_{\text{target}}^{i}}$
Then the domain weights are updated every $m$ steps as:
$w_{t}\leftarrow\frac{w_{t-m}[i]\cdot\exp(V_{t}[i])}{\sum_{i=1}^{k}w_{t-m}[i]\cdot\exp(V_{t}[i])}$
@@
This formulation dynamically adjusts domain weights based on model’s learning velocities in domains, ensuring an equitable comparison of learning efficiency across domains with varying initial and target losses. By doing so, we prioritize learning in domains where progress towards the target loss is most promising, thereby enhancing overall model performance. In summary, this approach dynamically rebalances the learning process across multiple domains, adjusting weights in response to observed rate of learning progress, and steering them towards achieving optimal and uniform performance improvements. The complete Velocitune algorithm is summarized in Algorithm 1.
@@
Require: Training data of $k$ domains $D_{1},D_{2},\cdots,D_{k}$, validation data $D^{\mathrm{val}}_{1},D^{\mathrm{val}}_{2},\cdots,D^{\mathrm{val}}_{k}$, initial data loading weights $\textbf{w}_{0}\in\mathbb{R}^{k}$ initialise as uniform distribution, upper bound loss $\ell_{\mathrm{U}}\in\mathbb{R}^{k}$, lower bound loss $\ell_{\mathrm{L}}\in\mathbb{R}^{k}$, LM loss $\mathcal{L_{T}}$ at training time $T$, evaluation per $m$ steps, model parameters $\theta$
@@
             $\Delta_{t}[i]\leftarrow\textnormal{{Clamp}}\left\{0,\frac{(\ell_{t}^{i}(\theta,S^{\prime}_{\text{eval},i})-\ell_{\mathrm{target}}[i])}{(\ell_{\mathrm{init}}[i]-\ell_{\mathrm{target}}[i])},1\right\}$triangleright Rescale training progress
@@
      Sample a batch of data $\mathcal{B}$ from $D_{1},D_{2},\cdots,D_{k}$ with proportion $w_{t}$;
@@
       $\alpha\leftarrow w\cdot\exp\left(\Delta\right)$triangleright Calculate the unnormalized weights
       $w\leftarrow\frac{\alpha}{\sum_{i}\alpha[i]}$ return $w$triangleright Renormalize and smoooth the data loading proportion
@@ 3 Experiment
Figure 2: Domain weights dynamic of Velocitune in training CodeLlama 7B on Reasoning.
@@
CodeLlama 20.9 18.8 21.6 30.50 43.20 36.80
CodeLlama-CPT 36.0 46.9 34.3 26.20 44.80 35.50
CodeLlama-Velocitune 37.3 56.2 35.9 (+1.6%) 34.10 44.40 39.3 (+3.8%)
Table 1: Performance of CodeLlama-7B, the baseline, and Velocitune on multiple math and code benchmarks.
@@ 3.1 Experimental setup
We compile the reasoning dataset based on Proof-Pile-2 (Azerbayev et al., 2023) which consists of math reasoning text in natural language, format language, and code. The dataset included three domains: Arxiv, AlgebraicStack, and OpenWebMath (Paster et al., 2023). Following the common practice to add replay data to prevent catastrophic forgetting, we add two more domains: general code and general language which are composed of the Github subset from SlimPajama (Soboleva et al., 2023) and a blend of Slimpajama except Github and Arxiv. This results in a training set spanning five domains, with 76% of the data in natural language and 24% in code. For system knowledge, we use a dataset SystemStack which is collected from three sources Arxiv, Blogs, and Stackoverflow, concentrating on computer system-related knowledge. The statistics of the two datasets are shown in Table 7.
@@
We trained the models using the standard next token prediction Negative Log-Likelihood (NLL) loss function. Three settings was compared, Velocitune, Dynamic Batch Loading (DBL) (Xia et al., 2023) which update weight using the distance between the evaluation loss and the target loss, and a baseline named CPT which is based on the number of tokens in each domain. For DBL, we also apply the predicted target loss. The total number of tokens processed during training was equivalent for the three methods to completing one full epoch using the training dataset. Detailed hyperparameters for the training process are summarized in Table LABEL:tab:training_details.
@@
To evaluate the math reasoning ability of models, we use math-lm-eval [^1] https://github.com/ZubinGou/math-evaluation-harness from ToRA (Gou et al., 2023) to evaluate accuracy on GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), Minerva (Lewkowycz et al., 2022),MMLU-STEM (Hendrycks et al., 2020), ASDiv (Miao et al., 2020), SVAMP (Patel et al., 2021), and MathQA (Amini et al., 2019). For coding ability, we use the evaluation kit from DeepSeek-Coder (Guo et al., 2024) to assess the pass@1 accuracy on HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021).
For evaluating command generation ability, we use two benchmarks CmdGen-NVIDIA and CmdGen-AMD from the CmdGen series, which are built to assess models’ ability to provide proper system command when asked by a related question. The two benchmarks, provided a combination of 1.5K instruction-tuning data and 205 and 192 test questions respectively, evaluate models output from six metrics: Similarity of command(CMD Sim): cosine similarity of CMD embeddings. Similarity of execution output(Output Sim): cosine similarity of output embeddings. Approximate accuracy: CMD Sim or Output Sim is larger than 0.9. The embeddings are generated by all-MiniLM-L6-v2222https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. Exact Match(EM): If model-generated commands are identical to the target commands. Success Ratio(SR): If the execution of the generated commands incur the same output from the system as the target command. Accuracy(Acc): The union of EM and SR. Example questions and commands are shown in Appendix A.

### v2

@@ Abstract
It is well-known that a diverse corpus is critical for training large language models, which are typically constructed from a mixture of various domains. In general, previous efforts resort to either sampling training data from different domains with static proportions or dynamically adjusting these proportions during training to optimise pretraining performance. However, few methods addressed the complexity of domain-adaptive continual pre-training. To fill this gap, we propose Velocitune, a novel framework that dynamically assesses learning velocity and adjusts data proportions accordingly, favouring slower learning domains while de-emphasising faster learning ones, which is guided by a scaling law to estimate the desired learning goal for each domain with a less associated cost. To evaluate the effectiveness of Velocitune, we conduct experiments on a dataset focused on reasoning tasks with CodeLlama, as well as on a corpus of system commands using Llama3 and Mistral. Velocitune achieves performance gains in both math and code reasoning tasks and command-line generation benchmarks. Further analysis reveals that key factors driving the effectiveness of Velocitune include target estimation and data ordering.
@@ 1 Introduction
Datasets used for pre-training language models (LMs) are typically composed of texts of various meta-attributes such as source and focus, referred to as different domains (Du et al., 2022; Azerbayev et al., 2023; Computer, 2023). The distinct characteristics of data from these varying domains, such as focus, quality, and quantity, affect the downstream performance of LMs differently (Rozière et al., 2024; Li et al., 2024). Consequently, numerous studies have explored the optimal combination of data from multiple domains to enhance LM performance.
@@
Llama3 (AI@Meta, 2024), GLaM (Du et al., 2022), and Lemma (Azerbayev et al., 2023) employ heuristic methods to iteratively test different ratios by training multiple proxy models and selecting the mixture that demonstrates the best downstream performance. However, these heuristic approaches demand costly large-scale experiments for effective exploration. As a result, recent research is focused on learning an optimal ratio by dynamically adjusting weights during proxy model training (Xie et al., 2024; Fan et al., 2023). For example, Doremi (Xie et al., 2024) implements a method in which a small reference model is initially trained, followed by training a small proxy model using group distributionally robust optimisation (Group DRO) (Sagawa et al., 2019a) to obtain optimised domain weights.
@@
Domain-adaptive [^1] Here, ”domain-adaptive” refers to optimising a model’s ability and knowledge in a specific field or area. It should not be confused with the concept of ”domain” of data as discussed in this paper. continual pre-training, while sharing some similarities with from-scratch training, presents unique challenges that limit the effectiveness of existing domain reweighting methods. Many existing methods utilise small proxy models to estimate optimal domain weights, which are subsequently transferred to a larger model (Xie et al., 2024; Fan et al., 2023; Azerbayev et al., 2023). However, this approach poses challenges in domain-adaptive continual pre-training, as smaller versions of the base model often do not exist, making it difficult to estimate weights from proxy models. Another challenge is how to leverage the learning status. Previous methods rely on the distance between the current loss and the target loss of the model for a domain  (Xie et al., 2024; Xia et al., 2023). However, this distance-based approach can result in overemphasis on specific domains, as domains with larger loss disparities may disproportionately influence learning. This can exacerbate imbalances across domains.
@@
To address these issues, we introduce a novel framework, Velocitune, centred on the concept of learning velocity, as illustrated in Figure 1. In contrast to previous approaches that leverage the distance between the current loss and the target loss, Velocitune more effectively captures how fast models learn in each domain by establishing the learning velocity. During training, domains exhibiting slower learning velocities are given increased weights, while those with faster velocities receive reduced weights, ensuring balanced learning progress.
To quantify learning velocity, it is crucial to determine both the model’s already learnt expertise and its desired learning goal for each domain. In the absence of a smaller proxy model, we leverage the Chinchilla scaling law Hoffmann et al. (2022), using the loss recorded on sub-sampled training data to cost-effectively predict the learning goal.
@@
We evaluated the performance of Velocitune in two settings: continual pre-training CodeLlama 7B (Rozière et al., 2024) on a math and coding reasoning dataset, as well as Llama3 and Mistral (Jiang et al., 2023) on a system command knowledge corpus. Velocitune demonstrates an average improvement of 1.6% in eight math tasks and 3.8% in two coding tasks compared to the baseline trained with default weights. In addition, Velocitune outperforms the baselines in Llama3, showing improvements of 4.9% and 3.1%, and in Mistral, with gains of 4.4% and 2.6% in the CmdGen-NVIDIA and CmdGen-AMD tasks, respectively. In addition, we conducted an in-depth ablation study to identify key factors contributing to the observed improvements. Our findings indicate that, beyond the contribution of the reweighted data ratios, the sequence of data ordering might also play a significant role in the effectiveness of Velocitune. The results show that incorporating the predicted target loss is critical for the effectiveness of the learning velocity. The contributions of this work can be summarised as follows.
@@
  • Introducing Velocitune, a novel framework for dynamically adjusting data ratios in continual pretraining. Velocitune estimates learning velocity to more precisely assess learning progress across domains and leverages scaling laws to optimise data allocation while minimising costs.
@@
  • Providing a detailed analysis revealing that reweighted data ratios, predicted target loss, and data ordering contribute to the effectiveness of Velocitune.
@@ 2 Domain-adaptive continual pre-training with Velocitune
In this section, we present a detailed explanation of Velocitune. Our approach focusses on adjusting domain weights based on learning velocity. To quantify learning velocity, it is essential to determine the model’s existing expertise in each domain, which can be represented by its initial evaluation losses. Additionally, we define a learning target as the expected loss that the model should achieve given a certain amount of training data. This target provides a reference for measuring progress and adjusting domain weights accordingly. During training, we periodically assess learning velocity, increasing the sampling weights of slower domains while reducing the weights of faster ones. This adjustment ensures balanced learning progress across domains. The methodology is detailed in § 2.1.
@@ 2.1 The Velocitune algorithm
Consider training a language model on a dataset $S$ consisting of $n$ distinct domains, denoted $D_{1},D_{2},\dots,D_{n}$. $S_{1},S_{2},\dots,S_{n}$ represent the subsets of data corresponding to each domain, where each $S_{i}$ is divided into a train and evaluation set, denoted as $S_{\text{train},i}$ and $S_{\text{eval},i}$, respectively. The weight of the domain $w\in\Delta^{n}$ represents the sampling weight assigned to the domains.
@@
We first evaluate the LM on each evaluation set $S_{\text{eval,},i}$ before training to obtain the initial loss $\ell_{\text{init}}$, providing an accurate measure of the model’s learnt expertise in each domain. The target loss $\ell_{\text{target}}$ for each domain is derived using the scaling law (Hoffmann et al., 2022). Specifically, we apply the Chinchilla scaling law by first defining the total dataset size and then training the model on subsets of training data using the default domain weights, which correspond to the ratio of tokens in each domain. Throughout the training process, multiple checkpoints are saved and the evaluation losses from these checkpoints are used to fit the scaling law parameters. Finally, we take the fitted function to predict the loss that the model could reach for using the entire dataset. Details of the implementation and analysis of the prediction errors are provided in Appendix B.
@@
After obtaining the initial loss and target loss in each domain, Velocitune iteratively updates domain weights, denoted by $w_{t}$, to minimise the re-weighted training loss effectively. The optimisation problem is formulated as follows, in two sequential steps:
(1) Minimise the weighted training loss:
$\min_{\theta}\sum_{t=1}^{T}\sum_{i=1}^{n}w_{t}[i]\cdot\ell_{t}^{i}(\theta)$ (1)
@@
where $T$ is the total training steps, $\ell_{t}^{i}(\theta)$ denotes the training loss at step $t$ for domain $D_{i}$ with model parameters $\theta$. The $w_{t}[i]$ refers to the weight assigned to domain $D_{i}$ at step $t$.
(2) Minimise the weighted sum of learning velocities:
$\min_{w\in\Delta^{n}}\sum_{i=1}^{n}w_{t}[i]\cdot V_{t}[i]$ (2)
@@
$V_{t}[i]$ is the learning velocity of domain $D_{i}$ at step $t$. The first step is to minimise the language modelling training loss given the domain weights. The second step focusses on dynamic adjustment so that the weight in the slowest learning velocity domain is maximised. Unlike Doremi (Xie et al., 2024), which updates domain weights based on the loss difference of each data point between a proxy model and a trained reference model, our method does not rely on a fully trained reference model. Consequently, the loss of each data point in the trained model is unobtainable. Instead, after a set number of training steps, we update the domain weights by first calculating the learning velocity, which reflects how quickly the model is learning in a given domain. The learning velocity is defined as:
$V_{t}[i]=\frac{\ell_{t}^{i}(\theta,S^{\prime}_{\text{eval},i})-\ell_{\text{target}}^{i}}{\ell_{\text{init}}^{i}-\ell_{\text{target}}^{i}}$ (3)
Here $S^{\prime}_{\text{eval}}$ is a subset of $S_{\text{eval}}$ to speed up velocity estimation during training and $\ell_{t}^{i}(\theta,S^{\prime}_{\text{eval},i})$ is the evaluation loss on $S^{\prime}_{\text{eval}}$ at step $t$ in domain $D_{i}$. Then the domain weights are updated every $m$ steps exponentially following Group DRO Sagawa et al. (2019b):
$w_{t}\leftarrow\frac{w_{t-m}[i]\cdot\exp(V_{t}[i])}{\sum_{i=1}^{n}w_{t-m}[i]\cdot\exp(V_{t}[i])}$ (4)
@@
Equation 4 adjusts the weights based on the learning velocity in each domain, ensuring an equitable comparison of the learning efficiency between domains with varying initial and target losses. By doing so, Velocitune prioritises learning in domains where progress towards the target loss is most promising, thereby enhancing overall model performance. In summary, this approach dynamically rebalances the learning process across multiple domains, adjusting weights in response to the observed rate of learning progress, and steering them towards achieving optimal and uniform performance improvements. The complete Velocitune algorithm is summarised in Algorithm 1.
@@
Require: Data of $n$ domains $S_{1},S_{2},\cdots,S_{n}$, initial data loading weights $w_{0}\in\mathbb{R}^{n}$ initialise as uniform distribution, initial loss $\ell_{\mathrm{init}}\in\mathbb{R}^{n}$, target loss $\ell_{\mathrm{target}}\in\mathbb{R}^{n}$, $\ell_{t}^{i}(\theta,S^{\prime}_{\text{eval},i})$ evaluation loss of the model at time $t$ , evaluation interval $m$, model parameters $\theta$.
@@
             $\Delta_{t}[i]\leftarrow\textnormal{{Clamp}}\left\{0,\frac{(\ell_{t}^{i}(\theta,S^{\prime}_{\text{eval},i})-\ell_{\mathrm{target}}[i])}{(\ell_{\mathrm{init}}[i]-\ell_{\mathrm{target}}[i])},1\right\}$triangleright Measure learning velocity
@@
      Sample a batch of data $\mathcal{B}$ from $S_{\text{train},1},S_{\text{train},2},\cdots,S_{\text{train},n}$ with proportion $w_{t}$;
@@
       $\alpha\leftarrow w\cdot\exp\left(\Delta\right)$triangleright Update unnormalised weights
       $w\leftarrow\frac{\alpha}{\sum^{n}_{i}\alpha[i]}$ return $w$triangleright Normalise domain weight
@@ 3 Experiment
Figure 2: Domain weights dynamic of Velocitune in training CodeLlama 7B on Reasoning. The vertical axis represents domain weights, while the horizontal axis denotes training steps.
Table 1: Performance of CodeLlama-7B, the baseline, and Velocitune on multiple math and code benchmarks. The highest average accuracy for math and code benchmarks is highlighted in bold.
@@
CodeLlama 20.9 18.8 21.6 30.5 43.2 36.8
CodeLlama-Baseline 36.0 46.9 34.3 26.2 44.8 35.5
CodeLlama-Velocitune 37.3 56.2 35.9 (+1.6%) 34.1 44.4 39.3 (+3.8%)
@@ 3.1 Experimental setup
We compile the Reasoning dataset based on Proof-Pile-2 (Azerbayev et al., 2023) which consists of math reasoning text in natural language, format language, and code. The dataset includes three domains: Arxiv, AlgebraicStack, and OpenWebMath (Paster et al., 2023). Following the common practice of adding replay data to prevent catastrophic forgetting, we add two more domains: general code and general language which are composed of the Github subset from SlimPajama (Soboleva et al., 2023) and a blend of Slimpajama except Github and Arxiv. This results in a training set spanning five domains, with 76% of the data in natural language and 24% in code. For system knowledge, we use a dataset SystemStack which is collected from three domains Arxiv, Blogs, and Stackoverflow, concentrating on computer system-related knowledge. The statistics of the two datasets are shown in Table 7.
@@
We trained the models using the Negative Log-Likelihood (NLL) loss. Three settings were compared, Velocitune, Dynamic Batch Loading (DBL) (Xia et al., 2023) which updates the weight using the distance between the evaluation loss and the target loss, and a baseline where sampling weights are determined by proportional token counts across domains during the continual pre-training. For DBL, we also apply the predicted target loss. The total number of tokens processed during training was equivalent for the three methods to completing one full epoch using the training dataset. Detailed hyperparameters for the training process are summarised in Appendix D LABEL:tab:training_details.
@@
To evaluate the math reasoning ability of models, we use math-lm-eval [^2] https://github.com/ZubinGou/math-evaluation-harness from ToRA (Gou et al., 2023) to evaluate the accuracy on GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), Minerva (Lewkowycz et al., 2022),MMLU-STEM (Hendrycks et al., 2020), ASDiv (Miao et al., 2020), SVAMP (Patel et al., 2021), and MathQA (Amini et al., 2019). For coding ability, we use the evaluation kit from DeepSeek-Coder (Guo et al., 2024) to assess the pass@1 accuracy on HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021).
For evaluating command generation ability, we use two benchmarks CmdGen-NVIDIA and CmdGen-AMD from the CmdGen series Lin et al. (2025), which are built to assess the ability of models to provide proper system command when asked a related question. The two benchmarks, which provided a combination of 1.5K instruction-tuning data and 205 and 192 test questions respectively, evaluate model output from six metrics: Similarity of Command(CMD Sim): Cosine similarity of embeddings of generated and target commands. Similarity of Execution Output(Output Sim): Cosine similarity of embeddings of system outputs. Approximate Accuracy(Approx Acc): A value of 1 is assigned if either CMD Sim or Output Sim exceeds 0.9; otherwise 0, the overall Approx Acc is aggregated over test cases. The embeddings are generated by all-MiniLM-L6-v2333https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. Exact Match (EM): The percentage of generated commands that are identical to the target commands. Success Ratio (SR): The percentage of generated commands that produce the same system output as the target command. Accuracy (Acc): The union of EM and SR, serving as the primary metric for the CMDGen task. Examples of CMDGen question and answer pair are shown in Appendix A.